OpenStudy (anonymous):

This is for integration properties of constants. I have lnx=ln(1+u^2)/2)+c when I take e to the both sides I get x=e^c (1+u^2)/2 the initial value problem is y(-1)=1 and the U becomes y/x I need to solve it for C which is the question how do I get C to the other side and solve it.

OpenStudy (anonymous):

The given constraint is y(-1) = 1... which means when x=-1, y=1... :-)

OpenStudy (anonymous):

You have to solve it for the value of C because of integration. The question is e^(c) what are the rules about moving the C around in e functions is e^c the same thing as c*e

hartnn (hartnn):

i suggest you find 'c' before taking e raised to both sides...

OpenStudy (anonymous):

I moved my equation around and I get c= x/sqrt(x^2+y^2/x^2)) I just am wondering what is the rules about constants. like if c*e/x^-1/2 is the same as e^c(x(1/2)

OpenStudy (anonymous):

why is e^c the same thing as c*e

hartnn (hartnn):

it isn't..

OpenStudy (anonymous):

Hey again... e^c and c^e are not same... And what I mean by *The given constraint is y(-1) = 1... which means when x=-1, y=1..* is that substitute x=-1, y=1, then you will get numerical value of 'c'... Hope this makes sense...

OpenStudy (anonymous):

I did do the substitution I just got confused on how I obtained c from the eqation lnx=ln(1+(y/x)^2)/2 + c or the lnx= sqrt(1+(y/x)^2)+c I get C= -1/sqrt(2) when I solve with initial values

hartnn (hartnn):

i would do it this way, x=e^c (1+u^2)/2 when x=-1, u= -1 so, e^c =-1 i think, you also got this ?? so, x=e^c (1+u^2)/2 --->x=-(1+u^2)/2

hartnn (hartnn):

no need to find c at all, use e^c=-1

OpenStudy (anonymous):

Thats what I did and I get e^c = -1/sqrt(2) so then i just put it into x=e^c(sqrt u^2+1) thats my answer ?

hartnn (hartnn):

also, c actually would be imaginary, e^c=-1 c= ln (-1) = i \(\pi\)

hartnn (hartnn):

how u got get e^c = -1/sqrt(2) ?

OpenStudy (anonymous):

This is where I'm getting confused... the same thing I got but it would give no real value of c... The only thing we can do from there is , use e^(ipi) = -1, that means taking log to -1 ... Otherwise if real c is asked, no such possible value... B/w I agree with hartnn... not getting the value you found...

OpenStudy (anonymous):

e^c x/sqrt of x^2+y^2/x^2 this value is what (y^2/x^2)+1 = then it is -1 =x and 1 = y so -1/sqrt((1+1/1)) -1/sqrt(2) the values of are all positive because of it being squared.

OpenStudy (anonymous):

and it is e^c=x/sqrt ((x^2+y^2)/x^2)

hartnn (hartnn):

ohh, its square root, then its correct...

OpenStudy (anonymous):

Thanks for your help again I think you helped me last night also. I got my seperation problem done is the ln(ln(u) anything I just left it as x=c-lnln(u)

hartnn (hartnn):

welcome :) yes. is your 2nd statement , a new problem ? i didn't get what u said...

OpenStudy (anonymous):

I asked you last night to about seperation of x(dy/dx)= yln(xy) I got it seperated in terms of v and x so dv/dx=-1lnv then integrated them to solve and i got lnln(v)=-x+c

hartnn (hartnn):

ohh! yes, now i remember.... actually i didn't solve it fully, so got confused , where did ln ln v came from...

OpenStudy (anonymous):

seperate the two gets dv/lnv =-1dx dv/u is ln(u) u=lnv so you get lnln(v)

hartnn (hartnn):

as far as i remember, u (or v) was xy and not in terms of ln....i hope you solved that correctly.

OpenStudy (anonymous):

dv/dx -v/x= (v/x)ln v this is variable separable, if you try it. is what you put but it simplifies into dv/dx= -1ln(v) v=xy and y=v/x

hartnn (hartnn):

ok, dv/dx = (v/x)(1+ln v) v(1+ln v) dv = dx/x now integrating (by parts for left, and ln x for right.)

hartnn (hartnn):

dv/dx= -1ln(v) is apparantely, not correct....

OpenStudy (anonymous):

uh oh I always forget h

OpenStudy (anonymous):

how to do stuff properly

hartnn (hartnn):

umm..practicing....